40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  ATF eLog  Not logged in ELOG logo
Message ID: 950     Entry time: Thu Aug 19 04:11:09 2010
Author: Dmass 
Type: Laser 
Category: Doubling 
Subject: PMC problem solved? 

I looked a little closer and found what I think are my problems. When I swept the cavity, the reflection dipped down to zero, so PMC is fine.

  • I was saturating my PDH sensing PD...I also noticed this when I swept the cavity.
  • I noticed that my locked PMC was oscillating at my modulation frequency (280 kHz)...so I added a few attenuators - final count: 13 - 6 - 90 = -83 dBm at laser PZT
  • I changed the sign of my feedback. Loop went to uber locked. Is it really possible that I was sitting on the wrong side (the tail) of my error signal?
  • Power through is now almost complete. I can no longer see the reflected beam on a sensitive card.
  • Measured: 1W transmission (with OPHIR high power meter), 38 mW reflection (with thorlabs meter).
ELOG V3.1.3-